Ifvthe ratio of fast readers to slow readers was 9 to 5, and there were 2,100 readers in all, how many were fast readers? (show your work)
Found 2 solutions by Nate, fractalier:
Answer by Nate(3500)

(9/(5 + 9))(2100) = fast readers
(9/14)(2100) = 1350 fast readers

Answer by fractalier(2101)

Let's call them f and s.
Thus the proportion we have is
f/s = 9/5
and cross-multiplying we have
5f = 9s
Also we know that f + s = 2100, or
s = 2100 - f
So we substitute this into the above and get
5f = 9(2100 - f)
5f = 18900 - 9f
14f = 18900
f = 1350 and
s = 750